# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v3

@@ 4.1 An Upper Bound on the Perplexity
To derive the first upper bound, we first provide an expression for the KL divergence between the end distributions $\bar{p}_{1}$ and $\bar{q}_{1}$ of two flows $\bar{p}_{t}$ and $\bar{q}_{t}$. To derive this expression, we extend the approaches of Opper & Sanguinetti (2007, Equation 3) and Haxholli et al. (2025) to DFM models.
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\tilde{w}_{t}^{i}(x^{i},x_{t})\log{\frac{\tilde{w}_{t}^{i}(x^{i},x_{t})}{\tilde{v}_{t}^{i}(x^{i},x_{t})}}+\tilde{v}_{t}^{i}(x^{i},x_{t})-\tilde{w}_{t}^{i}(x^{i},x_{t})\bigg)dt+D_{KL}(\bar{q}_{0}\|\bar{p}_{0}),$
@@
$D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\bigg)dt.$
A proof is provided in Appendix A.1. Motivated by the last result and Lou et al. (2024), we choose $\bar{p}_{t}(x)$ to be the learned approximation of flow $p_{t}$ in Equation (2) with the coupling $\pi(x_{0},x_{1})$, i.e., $\bar{p}_{t}(x)=p_{t}(x;\theta)$ and $v_{t}=u_{t}(x^{i},x_{t};\theta)$ . On the other hand, we choose $\bar{q}_{t}(x)$ to have the dynamics of $p_{t}$, but with the coupling $\bar{\pi}(x,y)=p_{0}(x)\delta_{x_{1}}(y)=\int\pi(x,z)dz\delta_{x_{1}}(y)$. Clearly, $\bar{q}_{0}(x)=p_{0}(x)$, $\bar{q}_{1}(x)=\delta_{x_{1}}(x)$ and $\bar{q}_{t}(x)=p_{t|1}(x|x_{1})$. We notice that since $\bar{q}_{0}(x)=p_{0}(x)$ and $\bar{p}_{0}(x)=p_{0}(x)$, then $D_{KL}(\bar{q}_{0}\|\bar{p}_{0})=0$. Furthermore $D_{KL}(\bar{q}_{1}(x)\|\bar{p}_{1}(x))=D_{KL}(\delta_{x_{1}}(x)\|p_{1}(x;\theta))=-\log{p_{1}(x_{1};\theta)}$. Thus, for such choices, $\log{p_{t}(x_{1};\theta)}$ is bounded from above by
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\bigg)dt.$ (13)
This bounds the negative-log-likelihood (NLL) for general DFM models. Shaul et al. (2025) concurrently obtained a similar result via an ELBO-based derivation. Taking expectations over $p_{1}(x_{1})$ on both sides gives a general bound on cross entropy $H(p_{1},p_{1}(\theta))$. For the dynamics from Equation (4),
$H(p_{1},p_{1}(\theta))$ $\leq\mathcal{B}:=\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}$
$\quad\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg)dt.$ (14)
A detailed derivation is provided in Appendix A.2. Hence $e^{\frac{\mathcal{B}}{L}}$ is a computable upper bound of the perplexity $e^{\frac{H(p_{1},p_{1}(\theta))}{L}}$ that can be used for training and evaluation (Algorithm 2 in Appendix B). Additionally, we provide an expression for the exact perplexity in Appendix A.5, but this cannot be used in practice as it requires knowing the learned probability ratios between neighbor states.
@@ 4.2 An Alternative Upper Bound on the Perplexity
$H(\bar{q}_{1})=H(\bar{q}_{0})+\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}}w_{t}^{i}(x_{t}^{i},x)\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}\bigg(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\bigg)dt,$ (15)
@@
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})+D_{KL}(\bar{q}_{0}\|\bar{p}_{0})+\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})$
$\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(\frac{\bar{q}^{i}_{t}(x|x_{0},x_{1})}{\bar{q}^{i}_{t}(x_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})\bigg)dt,$ (16)
@@ 5 Experiments
This section empirically validates our results. As proof of concept, we first show on small-vocabulary datasets that applying Theorem 3.1 effectively reduces jumps. Therein, we use the similarity metric $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$. We then confirm our bounds empirically, and in simple settings, estimate their tightness. In Section 5.3, we use $s(x^{i},x_{t}^{i})=\|e_{m}(x^{i})-e_{m}(x_{t}^{i})\|^{2}$, with learnable embeddings, where $e_{m}(k)$ denotes the $k$-th column of a learnable embedding matrix $E_{d\times V}$. Therein, we demonstrate that in realistic scenarios, applying Theorem 3.1 can reduce generation steps up to 8-fold to reach the same generative perplexity for BoW source distributions. Additionally, we show that multimask flows outperform masked flows, especially when combined with OT. Further, we calculate OT overhead, showing minibatch OT is practical. Finally, we conduct ablation studies on two key factors: OT batch size and OT solver. We use the POT implementation of Flamary et al. (2021) to compute optimal minibatch couplings throughout these experiments.
@@ 5.1 Proof of Concept Experiments
We trained a time-conditioned GPT-2 transformer with full attention on the Morse-code converted Shakespeare dataset, where non-convertible characters were left unchanged. The source sequence used was Bag-of-Words (BoW). A sample sequence from the BoW is constructed by sampling independently per position from the token frequencies in the training set. Training consisted of 100k iterations, character-level tokenization, sequence length 128, and batch size 64. We compared standard training with minibatch-OT. Minibatch OT increased training time by $0.3\%$ without affecting inference. We used Hamming distance and the Sinkhorn algorithm with entropy regularization parameter $\epsilon$. During inference with 1024 Euclidean steps, we counted token changes at each position across 3,000 generated sequences. Results appear in Table 1.
Since unstructured source sequences require modifying most tokens to generate structured data, there is a natural lower bound on required modifications. In Shakespeare Morse, the vocabulary contains three main tokens, each with probability $\sim$1/3. Thus, any given token has probability 2/3 of needing change, yielding an expected minimum of $128\times\frac{2}{3}=85.33$ jumps for sequence length 128. The standard method’s 85.47 jumps nearly matches this theoretical minimum, suggesting near-optimal performance. That OT reduces this to 74.84 is significant, demonstrating that OT-trained models generate samples closer to the source sequence while maintaining unbiased sampling when marginalizing across source sequences. Beyond jump counts, we report perplexity using the bound in Equation 14. For OT-trained models, we apply exact minibatch OT during testing to prevent sample repetition. Additionally, we report relative jumps, calculated as the ratio of each model’s jump count to that of the best performing model.
We also performed the same experiments on Shakespeare using a character-level tokenizer (see Appendix C.1). As discussed in Section 3, the increased vocabulary size makes Hamming distance less effective, necessitating the usage of the $L_{2}$ metric (Section 5.3) or other specialized measures.
@@
Model (L=128) Jumps Relative Jumps Perplexity
@@
With OT $\epsilon=0.01$ $\textbf{74.87}\pm 0.1$ 1 2.12
5.2 Utilizing the Bounds and Estimating Their Tightness
We test in practice the utility of our bounds as optimization targets and evaluation metrics. In Section 4.2, we mentioned that for masked DFM, both bounds coincide and simplify to $\int_{0}^{1}\frac{1}{1-t}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}-\delta_{m}(x_{t}^{i})\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}dt$. This matches the MD4 bound of Shi et al. (2024) for masked discrete diffusion (Appendix A.4). We denote models trained with this loss as DFM-S, those trained with the loss multiplied by (1t)(1-t) as DFM-N, and those trained with cross-entropy as DFM-O. Using the architecture from Section 5.1 with GPT2 tokenization, we trained on OpenWebText (OWT) (Gokaslan & Cohen, 2019) for 400K steps (batch size 512, sequence length 128). Testing on datasets from Lou et al. (2024), DFM-N performed best (Appendix C.2), so we compared DFM-N against SEDD and GPT-2 for longer sequences (L=1024). The performance gap between DFM-N and DFM-S mirrors that of CEDD and CEDDT in diffusion (Haxholli et al., 2025), further validating our bounds. Table 2 demonstrates that our bounds enable direct comparison with autoregressive models, with DFM-N and SEDD performing similarly.
@@ 5.2 Utilizing the Bounds and Estimating Their Tightness
5.3 Multi-masked Flows and Minibatch-OT on OpenWebText
To test minibatch-OT in practice, we trained a time-conditioned GPT-2-sized model with full attention for 400k iterations on OWT, using batch size 512 and sequence length 128. We compared: (1) a baseline (DFM-B) without OT, and (2) DFM-B-OT, that is, DFM-B trained using minibatch-OT with $L_{2}$ metric (Corollary 3.3). Both used the GPT-2 tokenizer (vocabulary size 50,257). The source distribution was BoW as in Section 5.1, but with OWT as the training set. The cross-entropy loss was used in both cases. After training, we generated 10,240 samples and evaluated quality using GPT2-large and Llama3.1 8B. OT significantly improved generative perplexity, reducing by 8-fold the generation steps (1024 to 128) needed to match the non-OT model’s score. Additionally, we measured the total transport cost in both dynamic and Kantorovich formulations for models trained with and without OT. The two formulations yielded similar values, and models trained with OT show lower transport costs as expected. See Appendix C.6 for details.
Table 3: Result differences between SOTA masked-flows, DFM-B, and DFM-MMLM with/without minibatch OT. GPT-2 Large was used as a judge. Asterisks denote the best results across all categories.
@@ 5.3 Multi-masked Flows and Minibatch-OT on OpenWebText
DFM-MMLM 536.50 288.38 204.77 170.61 155.45 143.48
DFM-MMLM-OT 525.83 283.10 199.55* 167.86* 153.51* 141.92*
Further, using the same experimental setting, we test multimasked flows with and without OT. Table 3 presents all generative perplexity results using GPT2 for evaluation, and the perplexity bound results are provided in Appendix C.3. In Appendix C.4, we provide the standard deviations, Llama evaluation results, and demonstrate through entropy scores that OT preserves the diversity. The Pearson correlation of perplexity and generative perplexity results is 0.96 (see Appendix C.7).

### v4

@@ 4.1 An Upper Bound on the Perplexity
For probability velocities $w$ and $v$, let $I^{i}_{t}(w,v,x^{i},x_{t})$ (or the shortened $I^{i}_{t}(w,v)$) denote
$w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t}).$
To derive the first upper bound, we first provide an expression for the KL divergence between the end distributions $\bar{p}_{1}$ and $\bar{q}_{1}$ of two flows $\bar{p}_{t}$ and $\bar{q}_{t}$. To derive this expression, we extend the approaches of Opper and Sanguinetti (2007, Equation 3) and Haxholli et al. (2025) to DFM models.
@@
$D_{KL}(\bar{q}_{1}\|\bar{p}_{1})=D_{KL}(\bar{q}_{0}\|\bar{p}_{0})$
$\int_{0}^{1}\sum_{i=1}^{L}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{x^{i}\neq x_{t}^{i}}\big(I^{i}_{t}(w,v)-I^{i}_{t}(\tilde{w},\tilde{v})\big)dt$ (13)
@@
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}I^{i}_{t}(w,v,x^{i},x_{t})dt+D_{KL}(\bar{q}_{0}\|\bar{p}_{0}).$
A proof is provided in Appendix A.1. Motivated by the last result and Lou et al. (2024), we choose $\bar{p}_{t}(x)$ to be the learned approximation of flow $p_{t}$ in Equation (2) with the coupling $\pi(x_{0},x_{1})$, i.e., $\bar{p}_{t}(x)=p_{t}(x;\theta)$ and $v_{t}=u(\theta)$ . On the other hand, we choose $\bar{q}_{t}(x)$ to have the dynamics of $p_{t}$, but with the coupling $\bar{\pi}(x,y)=p_{0}(x)\delta_{x_{1}}(y)=\int\pi(x,z)dz\delta_{x_{1}}(y)$. Clearly, $\bar{q}_{0}(x)=p_{0}(x)$, $\bar{q}_{1}(x)=\delta_{x_{1}}(x)$ and $\bar{q}_{t}(x)=p_{t|1}(x|x_{1})$. We notice that since $\bar{q}_{0}(x)=p_{0}(x)$ and $\bar{p}_{0}(x)=p_{0}(x)$, then $D_{KL}(\bar{q}_{0}\|\bar{p}_{0})=0$. Furthermore $D_{KL}(\bar{q}_{1}(x)\|\bar{p}_{1}(x))=D_{KL}(\delta_{x_{1}}(x)\|p_{1}(x;\theta))=-\log{p_{1}(x_{1};\theta)}$. Thus, for such choices, $\log{p_{t}(x_{1};\theta)}$ is bounded from above by
$\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}I^{i}_{t}(w_{t},u_{t}(\theta),x^{i},x_{t}).$ (14)
This bounds the negative-log-likelihood (NLL) for general DFM models. Shaul et al. (2025) concurrently and independently obtained a similar result via an ELBO-based derivation. Taking expectations over $p_{1}(x_{1})$ on both sides gives a general bound on cross entropy $H(p_{1},p_{1}(\theta))$. For the dynamics from Equation (4),
$H(p_{1},p_{1}(\theta))\leq\mathcal{B}$
$\int_{0}^{1}dt\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}$
$\bigg(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\bigg).$ (15)
A detailed derivation is provided in Appendix A.2. Hence $e^{\frac{\mathcal{B}}{L}}$ is a computable upper bound of the perplexity $e^{\frac{H(p_{1},p_{1}(\theta))}{L}}$ that can be used for training and evaluation (Algorithm 2 in Appendix B). in Appendix A.5, we provide an expression for the exact perplexity, but which in practice is computationally expensive and requires approximating the learned probability ratios between neighbor states.
@@ 4.2 An Alternative Upper Bound on the Perplexity
$H(\bar{q}_{1})=H(\bar{q}_{0})+$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}}w_{t}^{i}(x_{t}^{i},x)\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}\bigg(\log{\frac{\bar{q}_{t}(x)}{\bar{q}_{t}(x_{t})}}-1\bigg)dt$ (16)
@@
$H(\bar{q}_{1},\bar{p}_{1})\leq H(\bar{q}_{0})-\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\tilde{w}_{t}^{i}(x^{i},x_{t})+$
$\int_{0}^{1}\sum_{x_{0},x_{1}}\pi(x_{0},x_{1})\sum_{x_{t}}\bar{q}_{t}(x_{t}|x_{0},x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\bigg(v_{t}^{i}(x^{i},x_{t})$
$\frac{\bar{q}^{i}_{t}(x|x_{0},x_{1})}{\bar{q}^{i}_{t}(x_{t}|x_{0},x_{1})}\tilde{w}_{t}^{i}(x_{t}^{i},x)\log{\frac{\tilde{w}_{t}^{i}(x_{t}^{i},x)}{v_{t}^{i}(x^{i},x_{t})}}\bigg)dt+D_{KL}(\bar{q}_{0}\|\bar{p}_{0}),$ (17)
@@ 5 Experiments
This section empirically validates our results. As a proof of concept, we first show on small-vocabulary datasets that applying Theorem 3.1 effectively reduces jumps. Therein, we use the similarity metric $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$. We then confirm our bounds empirically, and in simple settings, estimate their tightness. In Section 5.3, we use the similarity measure $s(x^{i},x_{t}^{i})=\|e_{m}(x^{i})-e_{m}(x_{t}^{i})\|^{2}$, with learnable embeddings, where $e_{m}(k)$ denotes the $k$-th column of a learnable embedding matrix $E_{d\times V}$. There, we demonstrate that multimask flows outperform masked flows and that in realistic scenarios, applying Theorem 3.1 through minibatch-OT enables models to reach the same generative perplexity as their base counterparts with up to 8-times fewer steps. Further, we conduct ablation studies on three key factors: OT batch size, OT solver, and OT-EMA. Further, we conduct ablation studies on three key factors: OT batch size, the OT solver, and OT-EMA. Using an exponential moving average (EMA) to decouple model training from OT optimization allows us to further reduce the number of steps to 32 for the Bow experiments, yielding a 32× total reduction in steps. Finally, we calculate and analyze the OT overhead, showing that minibatch OT is practical and scales favorably. We use the POT implementation of Flamary et al. (2021) to compute optimal minibatch couplings throughout these experiments.
@@ 5.1 Proof of Concept Experiments
We trained a time-conditioned GPT-2 transformer with full attention on the Morse-code converted Shakespeare dataset, where non-convertible characters were left unchanged. The source sequence distribution used was Bag-of-Words (BoW). A sample sequence from the BoW is constructed by sampling independently per position from the token frequencies in the training set. Training consisted of 100k iterations, character-level tokenization, sequence length 128, and batch size 64. We compared standard training with minibatch-OT. Minibatch OT increased training time by $0.3\%$ without affecting inference. We used Hamming distance and the Sinkhorn algorithm with entropy regularization parameter $\epsilon$. During inference with 1024 Euclidean steps, we counted token changes at each position across 3,000 generated sequences. Results appear in Table 1.
@@
Model (L=128) Jumps RJ Perplexity
@@
With OT $\epsilon=0.01$ 74.87 $\pm$ 0.1 1 2.12
Since unstructured source sequences require modifying most tokens to generate structured data, there is a natural lower bound on required modifications. In Shakespeare Morse, the vocabulary contains three main tokens, each with probability $\sim$1/3. Thus, any given token has probability 2/3 of needing change, yielding an expected minimum of $128\times\frac{2}{3}=85.33$ jumps for sequence length 128. The standard method’s 85.47 jumps nearly matches this theoretical minimum, suggesting near-optimal performance. That OT reduces this to 74.84 is significant, demonstrating that OT-trained models generate samples closer to the source sequence while maintaining unbiased sampling when marginalizing across source sequences. Beyond jump counts, we report perplexity using the bound in Equation 15. For OT-trained models, we apply exact minibatch OT during testing to prevent sample repetition. Additionally, we report relative jumps (RJ), calculated as the ratio of each model’s jump count to that of the best performing model.
@@
We also performed the same experiments on Shakespeare using a character-level tokenizer (see Appendix C.1). As discussed in Section 3, the increased vocabulary size makes Hamming distance less effective, necessitating the usage of the $L_{2}$ metric (Section 5.3).
5.2 Utilizing the Bounds and Estimating Their Tightness
We test in practice the utility of our bounds as optimization targets and evaluation metrics. In Section 4.2, we mentioned that for masked DFM, both bounds coincide and simplify to $\int_{0}^{1}\frac{1}{1-t}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}$ $\big(-\delta_{m}(x_{t}^{i})\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}dt\big)$. This particular case matches the MD4 bound of Shi et al. (2024) for masked discrete diffusion (Appendix A.4). We denote models trained with this loss as DFM-S, those trained with the loss multiplied by (1t)(1-t) as DFM-N, and those trained with cross-entropy as DFM-O. Using the architecture from Section 5.1 with GPT2 tokenization, we trained on OpenWebText (OWT) (Gokaslan and Cohen, 2019) for 400K steps (batch size 512, sequence length 128). Testing on datasets from Lou et al. (2024), DFM-N performed best (Appendix C.2), so we compared DFM-N against SEDD and GPT-2 for longer sequences (L=1024). The performance gap between DFM-N and DFM-S mirrors that of CEDD and CEDDT in diffusion (Haxholli et al., 2025), further validating our bounds. Table 2 demonstrates that our bounds enable direct comparison with autoregressive models, as well as that DFM-N and SEDD perform similarly as expected.
Table 3: Result differences between SOTA masked-flows, DFM-B, and DFM-MMF with/without minibatch OT. GPT-2 Large was used as a judge. Asterisks denote the best results across all categories.
@@ 5.2 Utilizing the Bounds and Estimating Their Tightness
DFM-MMF 536.50 288.38 204.77 170.61 155.45 143.48
DFM-MMF-OT 525.83 283.10 199.55 167.86 153.51 141.92
DFM-MMF-Exact 518.86 281.39 199.68 168.12 153.51 141.46
DFM-MMF-Exact-EMA 480.61 259.44 187.68 156.62* 143.08* 132.55*
5.3 Minibatch-OT on OpenWebText
To test minibatch-OT in practice, we trained a time-conditioned GPT-2-sized model with full attention for 400k iterations on OWT, using batch size 512 and sequence length 128. We compared: (1) a baseline (DFM-B) without OT, and (2) DFM-B-OT, that is, DFM-B trained using minibatch-OT with $L_{2}$ metric (Corollary 3.3). Both used the GPT-2 tokenizer ($V_{d}=50,257$). The source distribution was BoW as in Section 5.1, but with OWT as the training set. The cross-entropy loss was used in both cases. After training, we generated 10,240 samples and evaluated quality using GPT2-large and Llama3.1 8B. OT significantly improved generative perplexity, reducing by 8-fold the generation steps (1024 to 128) needed to match the non-OT model’s score (Table 3). Additionally, we measured the total transport cost in both dynamic and Kantorovich formulations. Both formulations yielded similar, lower costs than models trained without minibatch-OT (Appendix C.6).
Further, using the same experimental setting, we test multimasked flows with and without OT. We choose the number of masks $V_{s}$ to match the number of data tokens $V_{d}=50,257$. Table 3 presents all generative perplexity results using GPT2 for evaluation, and the perplexity bound results are provided in Appendix C.3. In Appendix C.4, we provide the standard deviations, Llama evaluation results, and demonstrate through entropy scores that OT preserves the diversity. The Pearson correlation of perplexity bound values and generative perplexity results is 0.933 (Appendix C.7), showing very strong agreement.
5.4 Additional Ablation Experiments
We conduct ablation studies on two key factors: OT batch size and OT solver. Switching from Sinkhorn to exact solver does not yields significant changes in terms of generative perplexity (Table 3), but it improves bound values. Increasing minibatch OT size to 4096 (while maintaining flow batch size) slightly improves results at 32 and 64 steps, again with more pronounced improvements in bound values (Appendix C.8).
Moreover, flow matching with minibatch OT entails two coupled optimization processes: selecting the optimal coupling and training the flow model. Because the coupling can change from one iteration to the next, these shifts can destabilize the flow dynamics; at the same time, parameter updates alter embedding-based similarities, thereby changing the optimal coupling itself. To mitigate this feedback loop, we compute minibatch couplings using moving-average embeddings, decoupling the embeddings used for OT from those being updated by the network. This strategy (DFM-MMF-Exact-EMA and DFM-B-Exact-EMA) substantially improves generative perplexity (Table 3). Notably, DFM-B-Exact-EMA achieves better generative perplexity with only 32 steps than non-OT DFM-B with 1024 generation steps, without degrading entropy (Table 12; Appendix C.4).